OpenStudy (anonymous):

does anyone have the awnsers for algebra 1 unit 5 lesson 8 portfolio?

OpenStudy (anonymous):

is it algebra 1 b at CA

OpenStudy (anonymous):

I'm currently doing it I could use some help too I'm on task 1 c I have the answers to a and b

OpenStudy (anonymous):

its algebra 1 A at CA

OpenStudy (anonymous):

yeah

OpenStudy (anonymous):

ugh that sucks..i only take algebra 1 B..sorry :[[

OpenStudy (anonymous):

Oh darn uhm well omgineedhelp the rate of change for a is x(5)=y because 2(5)=10 x=2 x 5 and y is 10

OpenStudy (anonymous):

b I actually got off of a question on here that was from like a while ago 5x=3y=-2 The rate of change is Y=mx+b 3y= -5x-2 y= -5/3 x - 2/3 = -5/3 I used slope y intercept form to find it out.

OpenStudy (anonymous):

thank you!!!

OpenStudy (anonymous):

yeah I can find out the other answers just hang tight kay?

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

Im sorry I can't figure out c the only answer i got was the rate of change is 5 and that is more than likely wrong ill try task 2

OpenStudy (anonymous):

ok thank you for trying ill tell you if i find it

OpenStudy (anonymous):

thanks

OpenStudy (anonymous):

no problem

OpenStudy (anonymous):

found task 2 for abc and d let me get it down then ill give you the answers

OpenStudy (anonymous):

ok thanks

OpenStudy (anonymous):

y - 4 = (4-2) / (-1--5) (x--1) y-4 = (2/4)(x+1) y-4 = (1/2)x + 1/2 y = (1/2)x + (9/2)

OpenStudy (anonymous):

thats A and it used point slope

OpenStudy (anonymous):

Thank you where are you from? just out of curiosidy

OpenStudy (anonymous):

b.slope 2, y-intercept -4 y = mx + b y = 2x -4

OpenStudy (anonymous):

and Ohio

OpenStudy (anonymous):

b is obviously slope y intercept form y=mx+b

OpenStudy (anonymous):

(6,0) and (0,3) m = -3/6 = -1/2 y = (-1/2)x + 3 I used slope intercept form.

OpenStudy (anonymous):

passes through (1,2) with slope -5/3 2 = (-5/3)*1 + b b = 6/3+5/3 = 11/3 y (-5/3)x + (11/3)

OpenStudy (anonymous):

slope again